# Probing multi-particle bunching from intermittency analysis in relativistic heavy-ion collisions

### v1

@@ Abstract
It has been demonstrated that intermittency, a self-similar correlation regarding the size of the phase space volume, exhibits sensitivity to particle bunching within a system situated in the two-dimensional cell. Based on recent findings concerning fluctuations in charge particle density observed in central Au + Au collisions at various energies ranging from $\sqrt{s_{NN}}$ = 7.7 to 200 GeV at RHIC/STAR, it is proposed that analyzing intermittency in relativistic heavy-ion collisions could serve as a method to investigate density fluctuations linked to correlation phenomena. Our understanding of particle production mechanisms may be enriched by using intermittency analysis to gain new information contained in them that was previously unavailable.
@@ I Introduction
The major goal of the intermittency analysis in relativistic heavy ion collision programs is to explore the phase diagram of quantum chromodynamic (QCD) matter. The critical end point (CEP) is a key feature of the QCD phase diagram, representing the point where the first-order phase transition boundary terminates [1, 2]. Many efforts have been made to search for the possible CEP in heavy-ion collisions [3, 4].
@@
where $M^{D}$ is the number of cells in D-dimensional phase space and $n_{i}$ is the measured multiplicity of a given event in the $i$-th cell. The angle bracket denotes an average over the events. The intermittency appears as a power-law (scaling) behavior of scaled factorial moments [14, 20, 16, 23]. If the system features density fluctuations, scaled factorial moments will obey a power-law behavior of $F_{q}\left(M\right)\propto\left(M^{D}\right)^{\phi_{q}}$, $M\gg 1$, where $\phi_{q}$ is called the intermittency index quantifying the strength of intermittency [14, 15, 16, 18, 21].
However, the consideration of ”intermittency” as a genuinely novel phenomenon remained less apparent. Relation between Hanbury-Brown-Twiss effect and intermitency phenomenon was discussed and advocate that they are perfectly compatible [25, 26]. Has been shown that the multi-particle Bose-Einstein correlations leads to the power-law like behavior of scaled factorial moments. In Ref. [27], diverse datasets were analyzed, leading to the conclusion that intermittency results from Bose-Einstein correlations, alongside a mechanism driving power-law behavior. Conversely, in Ref. [28], the authors contend that the observed intermittent-like patterns in moments of multiplicity distributions are attributable to the quantum statistical properties of the particle-emitting system, rather than serving as direct evidence of intermittency.
@@
Recently the STAR Collaboration published extensive data on $\Delta F_{q}\left(M\right)$ in two dimensional ($D=2$) transverse momentum plane ($p_{Tx},p_{Ty}$) from central Au+Au collisions at $\sqrt{s_{NN}}$ = 7.7, 11.5, 19.6, 27, 39, 62.4 and 200 GeV [29]. In the next sections, based on the STAR experiment data, we provide a discussion concerning multi-particle bunching and suggest that measuring the intermittency in relativistic heavy-ion collisions could serve as a means to investigate density fluctuations within a highly constrained number of cells. This observation will be our main point for further discussion and calculations described in next sections.
@@ V Possible origin of power-law behavior
Intermittency appears as a power-law behavior of scaled factorial moments, $\Delta F_{q}\propto(M^{2})^{\Phi_{q}}$. For a scenario in which a bunch of particles appears only in one cell (or in a limiting number of cells), following Eq. (5), we expect the intermittency index $\Phi_{q}=q-1$, for multiplicity $m$ of particles in a bunch being independent on the number of cells $M^{2}$. In reality, the division procedure of phase space changes multiplicity $m$ in a cell. Number of particles in bunch of effective space size $\delta$ in cell of size $\Delta\left(M\right)=\Delta\left(M=1\right)/M$ is expected to be $m=m(\delta/\Delta(M))=m(M)$.
@@
Numerical results of the $m\left(M^{2}\right)$ dependence are shown in Fig. 8. For large $M^{2}$ we observe power-law dependence:
@@
$\Delta F_{q}\left(M^{2}\right)\propto m^{q}\left(M^{2}\right)\left(M^{2}\right)^{q-1}=\left(M^{2}\right)^{q\left(1-\kappa\right)-1}$ (17)
The intermittency index $\Phi_{q}=q\left(1-\kappa\right)-1$ is the result of the division procedure, which leads to power-law dependence $m\left(M^{2}\right)$ given by Eq. (16). It is known [35] that, if function $m\left(M^{2}\right)$ follows a simple power law, $m\left(M^{2}\right)\propto\left(M^{2}\right)^{-\kappa}$ then it is scale invariant and

### v2

@@ Abstract
It has been demonstrated that factorial moments analysis in dependence from the size of phase-space cells (when the latter is decreased but is still considerably large), exhibits sensitivity to particle bunching within a system situated in the two-dimensional cell. Based on recent findings concerning fluctuations in charge particle density observed in central Au + Au collisions at various energies ranging from $\sqrt{s_{NN}}$ = 7.7 to 200 GeV at RHIC/STAR, it is proposed that analyzing factorial moments in relativistic heavy-ion collisions could serve as a method to investigate density fluctuations linked to correlation phenomena. Our understanding of particle production mechanisms may be enriched by using factorial moments analysis to gain new information contained in them that was previously unavailable.
@@ I Introduction
The major goal of the intermittency (a self-similar correlation regarding the size of the phase-space volume) analysis in relativistic heavy ion collision programs is to explore the phase diagram of quantum chromodynamic (QCD) matter. The critical end point (CEP) is a key feature of the QCD phase diagram, representing the point where the first-order phase transition boundary terminates [1, 2]. Many efforts have been made to search for the possible CEP in heavy-ion collisions [3, 4].
@@
where $M^{D}$ is the number of cells in D-dimensional phase space and $n_{i}$ is the measured multiplicity of a given event in the $i$-th cell. The angle bracket denotes an average over the events. The intermittency appears as a power-law (scaling) behavior of scaled factorial moments, and factorial moments increase when the size of the phase-space cells goes to zero [^1] Frequently in multiparticle production, a system is said to be ”intermittent” if a power-law relations of factorial moments exist [25]. Usually one observes an increase of factorial moments with decreasing size of phase-space cells but this happens for quite large cells and not in the limit where the size of phase-space cells goes to zero. However, for smaller cells the factorial moments saturate and no intermittency effect can be observed. [14, 20, 16, 23]. If the system features density fluctuations, scaled factorial moments will obey a power-law behavior of $F_{q}\left(M\right)\propto\left(M^{D}\right)^{\phi_{q}}$, $M\gg 1$, where $\phi_{q}$ is called the intermittency index quantifying the strength of intermittency [14, 15, 16, 18, 21].
However, the consideration of ”intermittency” as a genuinely novel phenomenon remained less apparent. Relation between Hanbury-Brown-Twiss effect and intermitency-like phenomenon was discussed and advocate that they are perfectly compatible [26, 27]. Has been shown that the multi-particle Bose-Einstein correlations leads to the power-law like behavior of scaled factorial moments. In Ref. [28], diverse data sets were analyzed, leading to the conclusion that intermittency results from Bose-Einstein correlations, alongside a mechanism driving power-law behavior. Conversely, in Ref. [29], the authors contend that the observed intermittent-like patterns in moments of multiplicity distributions are attributable to the quantum statistical properties of the particle-emitting system, rather than serving as direct evidence of intermittency.
@@
Recently the STAR Collaboration published extensive data on $\Delta F_{q}\left(M\right)$ in two dimensional ($D=2$) transverse momentum plane ($p_{Tx},p_{Ty}$) from central Au+Au collisions at $\sqrt{s_{NN}}$ = 7.7, 11.5, 19.6, 27, 39, 62.4 and 200 GeV [25]. In the next sections, based on the STAR experiment data, we provide a discussion concerning multi-particle bunching and suggest that measuring the scaled factorial moments in relativistic heavy-ion collisions could serve as a means to investigate density fluctuations within a highly constrained number of cells. This observation will be our main point for further discussion and calculations described in next sections.
@@ V Possible origin of power-law behavior
For a scenario in which a bunch of particles appears only in one cell (or in a limiting number of cells), following Eq. (5), we expect a power-law behavior of scaled factorial moments, $\Delta F_{q}\propto(M^{2})^{\Phi_{q}}$ with the power index $\Phi_{q}=q-1$, for multiplicity $m$ of particles in a bunch being independent on the number of cells $M^{2}$. In reality, the division procedure of phase space changes multiplicity $m$ in a cell. Number of particles in bunch of effective space size $\delta$ in cell of size $\Delta\left(M\right)=\Delta\left(M=1\right)/M$ is expected to be $m=m(\delta/\Delta(M))=m(M)$.
@@
Numerical results of the $m\left(M^{2}\right)$ dependence are shown in Fig. 8. For large $M^{2}$ we observe power-law dependence [^3] The procedure of division of particles in a given bunch occurs as a statistical process at a deformed exponential rate. This means that the number of particles in a bunch that are likely to be removed from a cell within a given cell number range is proportional to the number of particles in a cell, $dm/d\left(M^{2}\right)=-m/\sigma^{2}$. In the considered cell division procedure, the division rate $\sigma^{2}=\sigma^{2}\left(M^{2}\right)$ depends on the numbers of cells, $M^{2}$ (for a large number of cells, the size of a single cell is small and there is more probable that the particles in a bunch are separated between different cells). For $\sigma^{2}\left(M^{2}\right)=\sigma_{0}^{2}+M^{2}/\kappa$ one has solution in the form of kappa-deformed exponential function (exponential function appears in the limit $\kappa\rightarrow\infty$), $m\left(M^{2}\right)\propto[1+M^{2}/\left(\sigma_{0}^{2}\kappa\right)]^{-\kappa}$. For large number of cells, $M^{2}\gg\sigma_{0}^{2}\kappa$, one has scale-free power-law dependence given by Eq. (16).:
@@
$\Delta F_{q}\left(M^{2}\right)\propto m^{q}\left(M^{2}\right)\left(M^{2}\right)^{q-1}=\left(M^{2}\right)^{q\left(1-\kappa\right)-1}.$ (17)
The power index $\Phi_{q}=q\left(1-\kappa\right)-1$ is the result of the division procedure, which leads to power-law dependence $m\left(M^{2}\right)$ given by Eq. (16). It is known [35] that, if function $m\left(M^{2}\right)$ follows a simple power law, $m\left(M^{2}\right)\propto\left(M^{2}\right)^{-\kappa}$ then it is scale invariant and